# Strong Haken via Thin Position

### v1

@@ 2.1. A word on notation
This is a by now standard sweepout argument so we simply sketch the proof. For more details, see (for example) [RS]. Let $\Gamma_{-}$ and $\Gamma_{+}$ be spines for the compressionbodies on either side of $M$, transverse to $\mathcal{S}$. Their complement is a product which allows us to define a continuous function $\phi\mskip 0.5mu\colon\thinspace M\to I$. We may perturb it so $\phi|_{\mathcal{C}}$ is Morse. Note that $\phi^{-1}(0)=\Gamma_{-}$, $\phi^{-1}(1)=\Gamma_{+}$ and $H_{t}=\phi^{-1}(t)$ is isotopic to $H$ for any $t\in I\setminus\partial I$. The critical values of $\phi$ divide $I$ into open sub-intervals. Label a subinterval with -- (resp. $+$) if for some (equivalently, any) $t$ in the subinterval has the property that there is an essential loop in $H_{t}\cap\mathcal{S}$ that is essential in $\phi^{-1}(t)$ and which bounds a disc below $H$ (resp. above $H$). As $H$ is strongly irreducible no interval is labelled with opposite signs. Since both $H$ and $\mathcal{C}$ are orientable, it is also the case that no two adjacent intervals have opposite sign.
Consider either the lowest interval or the highest interval. Let $t$ be in the interval and let $C$ be the compressionbody below or above $H_{t}$ respectively. We think of $H_{t}$ as being the frontier of a regular neighborhood of the spine. Let $S_{0}\subset\mathcal{S}$ be a component. Each intersection of $S_{0}$ with $H$ is either a loop in $C$ parallel to $\partial S_{0}$ or a meridian of an edge of the spine. A meridian intersection bounds an s-disc for $H$ in $C$. If some such disc is a compressing disc, the interval is labelled -- (if it is the lowest interval) or $+$ (if it is the highest interval). If not, $H_{t}$ is the surface isotopic to $H$ satisfying the conclusions of the lemma.
Assume, therefore, that both sides of $H$ admit compressing discs, so the lowest interval is labelled -- and the highest is labelled $+$. We conclude that there must be a subinterval with no label. Let $t$ lie in that subinterval. If $H_{t}\cap\mathcal{S}=\varnothing$, we are done, so suppose that $\gamma\subset H_{t}\cap\mathcal{S}$ is a component. As $\mathcal{S}$ is the union of discs and spheres, we may assume that it is innermost in $\mathcal{S}$. Let $D\subset\mathcal{S}$ be the innermost disc it bounds. Since the interval containing $t$ has no label, $\partial D$ is inessential on $H$. If $D$ is not a semi-compressing disc for $H$, we can isotope $H$ so as to remove $\partial D$ from $H\cap\mathcal{H}$. ∎
@@ 4. Heegaard surfaces through thick and thin
Following [HS], we reinterpret [ST] in terms of surfaces, rather than handle structures. Suppose that $\mathcal{H}$ is an oriented generalized Heegaard surface and that $H$ is an weakly reducible thick surface. Let $D_{-}$ and $D_{+}$ be unions of pairwise disjoint s-discs for $H$, with $D_{-}$ below $H$ and $D_{+}$ above $H$. Let $H_{\pm}=H|_{D_{\pm}}$ and let $F=H|_{D_{-}\cup D_{+}}$. A small isotopy of $H_{+}$ above $H$ and $H_{-}$ below $H$ makes $H_{-}$, $H_{+}$, and $F$ mutually disjoint. Let $\mathcal{J}^{+}=(\mathcal{H}\setminus H)\cup H_{-}\cup H_{+}$ and let $\mathcal{J}^{-}=\mathcal{H}^{-}\cup F$. Then we say that $\mathcal{J}$ is obtained by an untelescoping $\mathcal{H}$ and that $D_{-}$ and $D_{+}$ are an s-weak reducing pair. It is a weak reducing pair if both $D_{-}$ and $D_{+}$ contain a compressing disc and a (s/2)-weak reducing pair if at least one of $D_{-}$ or $D_{+}$ contain a compressing disc. See Figure 3.
@@
On the other hand, suppose $\mathcal{H}$ contains a thick surface $H$ and a thin surface $F$ such that $H\cup F$ lie in the boundary of a product compressionbody component $C$ of $M\setminus\mathcal{H}$. We say that $\mathcal{J}=\mathcal{H}\setminus(\mathcal{H}\cup F)$ is obtained by consolidating $\mathcal{H}$.
@@
    untelescoping using an s-weak reducing pair $D_{-}$, $D_{+}$, such that one of $D_{\pm}$ contains a compressing disc and a positive genus component of $H_{\mp}=H|_{D_{\mp}}$ can be consolidated with a component of $\mathcal{H}^{-}$; we also perform that consolidation.
@@
If $\alpha,\beta$ are two such complexity sequences, we say that $\alpha<\beta$ if and only if there exists $k\geq 0$ such that the first $k$ terms of $\alpha$ and $\beta$ coincide and either $\alpha$ does not have a $(k+1)$st term, but $\beta$ does (possibly it is 0) or the $(k+1)$st term of $\beta$ is strictly greater than the $(k+1)$st term of $\alpha$. Observe that complexities are well-ordered.
Consolidation removes a term from a complexity sequence and so decreases it. A type I untelescoping replaces an entry in a complexity sequence with some number (at least two) of other terms, each strictly smaller than the original. Thus, it also decreases complexity. Finally, consider a type II untelescoping and consolidation. Assume the untelescoping using an s-weak reducing pair $D_{-}$, $D_{+}$ such that $D_{-}$ contains only semi-compressing discs and $D_{+}$ contains at least one compressing disc. Let $H\subset\mathcal{H}^{+}$ contain $\partial D_{-}\cup\partial D_{+}$. Since $H$ admits a compressing disc, $c(H)>0$. Then if $H^{\prime}$ is a component of $H_{+}=H|_{D_{+}}$, we have $c(H^{\prime})<c(H)$. If $H^{\prime}$ is a component of $H_{-}=H|_{D_{-}}$, either $c(H^{\prime})=c(H)$ or $c(H^{\prime})=0$. In the former case, we immediately consolidate the component $H^{\prime}$, and so the type II untelescoping and consolidation also strictly decreases complexity. ∎
@@
Suppose that $\mathcal{H}$ is locally thin oriented generalized Heegaard surface for $M$. Then the following hold:
@@
If $\mathcal{H}$ is locally thin, we cannot guarantee that each component of $\mathcal{H}$ is essential; there may be components that are 2-spheres bounding 3-balls or components that are $\partial$-parallel. Such thin surfaces will not interfere in what follows. As in [ST, HS, TT1], if we allowed certain destabilization operations, we could continue thinning and ensure the thin surfaces were essential.
@@
Assume that $M$ is a reducible manifold such that no component of $\partial M$ is a sphere. If $\mathcal{H}$ is a locally thin generalized Heegaard surface for $M$, then $\mathcal{H}^{-}$ contains a collection of essential spheres $P$ such that each component of $M|_{P}$ is irreducible. In particular, if $M$ contains a non-separating sphere, then so does $\mathcal{H}^{-}$.
@@
Let $P\subset\mathcal{H}^{-}$ be the union of all the spheres in $\mathcal{H}^{-}$ that are essential in $M$. Let $M_{0}$ be a component of $M|_{P}$ and let $\mathcal{H}_{0}$ be those components of $\mathcal{H}$ contained in the interior of $M_{0}$. Using the fact that $\mathcal{H}$ cannot be consolidated and that the existence of spherical components in $\partial_{-}C$ for a compressionbody $C$, does not affect which curves in $\partial_{+}C$ bound compressing discs in $C$, we see that $\mathcal{H}_{0}$ is also thin.
@@
Since the previous paragraph modifies the sphere $Q$ by surgery, and since every sphere in a compressionbody is separating, if $Q$ were non-separating, we see that $P$ must also have contained a nonseparating sphere. ∎
@@ 5. Amalgamation
Suppose that $M$ is a 3-manifold containing an transversally oriented closed surface $F\subset M$ (possibly disconnected) separating $M$ into (possibly disconnected) 3-manifolds $M_{-}$ and $M_{+}$, such that the transverse orientation on $F$ points into each component of $M_{+}$ and out of each component of $M_{-}$ and so that each component of $M_{-}$ and $M_{+}$ is incident to a component of $F$. Suppose that $H_{\pm}$ is the union of oriented Heegaard surfaces for $M_{\pm}$ such that the orientation on $H_{+}$ (resp. $H_{-}$) points away from (resp. towards) $F$. We can amalgamate the surfaces $H_{-}$ and $H_{+}$ across $F$ as follows. See Figure 4.
@@
For notational convenience, Let $U_{\pm},V_{\pm}$ be the union of the compressionbodies on either side of $H_{\pm}$ with $F\subset U_{-}\cap V_{+}$. Choose a complete collection of discs for each of $U_{-}$ and $V_{+}$. When we $\partial$-reduce $U_{-}\cup V_{+}$ along those discs, we obtain product compressionbodies and 3-balls. The union of some of those product compressionbodies is $\Pi=F\times[-1,1]$ where $F\times[-1,0]\subset V_{+}$ and $F\times[0,1]\subset U_{-}$. In both cases, consider $F$ as $F\times\{0\}$. The other product compressionbodies contain $\partial_{-}V_{+}\setminus F$ and $\partial_{-}U_{-}\setminus F$. Let $\Delta_{-}\subset U_{-}$ be those discs in the complete collection that leave scars on $F\times\{1\}$. A small isotopy of the product structure on $\Pi$ ensures that the projection of the scars from $\Delta_{-}$ on $F\times\{1\}$ to $F\times\{-1\}$ completely misses the scars in $F\times\{-1\}$ arising from the complete collection of discs in $V_{+}$. Let $A\subset\Pi$ be the solid tubes corresponding to the projections. Let $X_{-}$ be the union of $V_{-}$ with $A$ and with the components of $U_{-}|_{\Delta_{-}}$ not containing $F$. Let $X_{+}$ be the closure of the complement of $X_{-}$ and let $H=X_{-}\cap X_{+}$. See Figure 4.
@@
Suppose that $M_{-}$ and $M_{+}$ are 3-manifolds such that $F=\partial M_{1}\cap\partial M_{2}$ is nonempty and oriented so that its orientation points into $M_{+}$ and out of $M_{-}$ Let $H_{-}$ and $H_{+}$ be oriented Heegaard surfaces for $M_{-}$ and $M_{+}$ respectively such that the orientation for $H_{+}$ points away from $F$ and the orientation for $H_{-}$ points toward $F$. Let $H$ be the Heegaard surface resulting from amalgamating $H_{+}$ and $H_{-}$ across $F$. Then:
@@
    If $F$ is a sphere, then $H$ is isotopic to the connected sum of $H_{-}$ and $H_{+}$ as Heegaard surfaces for the 3-manifolds $M_{-}$ and $M_{+}$ obtained by capping off $F$ with 3-balls.
@@
A height function on an oriented generalized Heegaard surface $\mathcal{H}$ is a function $f\mskip 0.5mu\colon\thinspace\mathcal{H}\to\mathbb{N}$ which is constant on each component of $\mathcal{H}$ and satisfying:
@@
If $\mathcal{J}$ is obtained from $\mathcal{H}$ by an amalgamation determined by a height function $f$, then $\mathcal{J}$ is acyclic and $f$ naturally induces a height function for $\mathcal{J}$.
@@
Use the notation of Definition 5.8. Without loss of generality assume that $C$ is below $H$. This implies that each $C^{\prime}_{i}$ is above $H_{i}$. Let $C_{i}$ be the compressionbody below $H_{i}$ and $C^{\prime}$ the compressionbody above $H$. Let $H^{\prime}$ be the surface resulting from the amalgamation. Let $U$ and $V$ be the compressionbodies above and below $H^{\prime}$ respectively. Let $\mu=f(H_{i})$ and note this is the same for all $i$.
@@
$f^{\prime}(F)=f(F)<f(H_{i})=f^{\prime}(H^{\prime}).$
@@
Figure 5. An example of an amalgamation respecting a height function. The numbers indicate the heights of the thick and thin surfaces. The top thick surfaces is amalgamated with the next two thick surfaces on in the left two columns but not the next thick surface in the rightmost column. Those are the two thick surfaces that are amalgamated with the topmost thick surface since they have the largest heights of the thick surfaces immediately adjacent to the topmost thick surface. The fact that the top right thin surface has a higher height is irrelevant.
@@
Let $\mathcal{H}$ be an oriented generalized Heegaard surface with height function $f$ and let $s$ be a sequence of amalgamations consistent with the induced height functions. For any compressionbody $C\subset M\setminus\mathcal{H}$ with $C$ below $\partial_{+}C$, if $F$ is a component of $\partial_{-}C$ with $f(F)<f(\partial_{+}C)$, add $(f(\partial_{+}C)-f(F)+1)/2$ parallel copies of $F$ to $\mathcal{H}$. One component in each copy is a thick surface and one is a thin surface and, together with $F$ and $\partial_{+}C$, the thick and thin surfaces alternate. Observe that we can restore $\mathcal{H}$ by consolidating the parallel copies we just inserted. We can extend both the height function $f$ and the sequence of amalgamations to incorporate these new copies. As amalgamating across a product region is the same as consolidation, this does not affect the result. Consequently, the result of amalgamation consistent with a height function, is independent of the height function, depending only on $\mathcal{H}$ and on the order of thick surfaces chosen in $s$. We now show how to eliminate even this latter dependency.
Let $\mathcal{H}^{\prime}$ be the new generalized Heegaard surface after one of the amalgamations. Its dual digraph is the result of subdividing edges of the dual digraph of $\mathcal{H}$. Let $H_{2i}=f^{-1}(2i)$ and $F_{2i-1}=f^{-1}(2i-1)$ for each $i\in\mathbb{N}$. Observe that if $C$ is a compressionbody below $\partial_{+}(C)$, then
@@
We may as well assume, therefore, that for each term of the amalgamation sequence $s$, there is an $i\in\mathbb{N}$ such that the thick surfaces at height $i$ are amalgamated with the thick surfaces at height $i-1$. It then follows from [Lackenby, Prop. 3.1] that $H$ is, up to isotopy, independent of $s$ and of the choice of disc sets used in each amalgamation. ∎
@@
The case when $\partial M$ contains spheres follows relatively easily from the case when it does not; for simplicity, assume that $\partial M$ does not contain spheres. For a generalized Heegaard surface $\mathcal{H}$, let $\operatorname{net}\chi(\mathcal{H})=-\chi(\mathcal{H}^{+})+\chi(\mathcal{H}^{-})$. It is easy to verify that if $\mathcal{H}$ thins to $\mathcal{J}$ or if $\mathcal{J}$ amalgamates to $\mathcal{H}$, then $\operatorname{net}\chi(\mathcal{H})=\operatorname{net}\chi(\mathcal{J})$. We treat the case of separating and non-separating spheres separately; in each case, proving two inequalities.
@@
Suppose that $M$ contains a non-separating sphere $P$. Let $H$ be a Heegaard surface for $M|_{P}$. One one side of $H$, choose a sphere $F$ bounding a 3-ball $B$. Inside $B$ remove two open 3-balls $B_{1},B_{2}$, creating two spherical boundary components $F_{1},F_{2}$. Let $W=S^{2}\times I$ and let $H^{\prime}\subset W$ be a sphere bounding a 3-ball in $W$. Let $M^{\prime}$ be the result of gluing the components of $\partial W$ to $F_{1}\cup F_{2}$. Then it is possible to assign orientations to $\mathcal{H}=H\cup F_{1}\cup F_{2}\cup H^{\prime}$ so that it is an oriented, acyclic generalized Heegaard surface with $H,H^{\prime}\subset\mathcal{H}^{+}$ and $F_{1},F_{2}\subset\mathcal{H}^{-}$. Notice that $M^{\prime}$ is homeomorphic to $M$ and that

### v2

@@ 2.1. A word on notation
This is a by now standard sweepout argument so we simply sketch the proof. For more details, see (for example) [RS]. Let $\Gamma_{-}$ and $\Gamma_{+}$ be spines for the compressionbodies on either side of $H$, transverse to $\mathcal{S}$. Their complement is a product which allows us to define a continuous function $\phi\mskip 0.5mu\colon\thinspace M\to I$. We may perturb it so $\phi|_{\mathcal{S}}$ is Morse. Note that $\phi^{-1}(0)=\Gamma_{-}$, $\phi^{-1}(1)=\Gamma_{+}$ and $H_{t}=\phi^{-1}(t)$ is isotopic to $H$ for any $t\in I\setminus\partial I$. The critical values of $\phi$ divide $I$ into open sub-intervals. Label a subinterval with - (resp. ++) if some (equivalently, any) $t$ in the subinterval has the property that there is a loop in $H_{t}\cap\mathcal{S}$ that is essential in $H_{t}$ and which bounds a disc below $H_{t}$ (resp. above $H_{t}$). As $H$ (and, hence, $H_{t}$) is strongly irreducible no interval is labelled with opposite signs. Since both $H$ and $\mathcal{S}$ are orientable, it is also the case that no two adjacent intervals have opposite sign.
Consider either the lowest interval or the highest interval. Let $t$ be in the interval and let $C$ be the compressionbody below or above $H_{t}$ respectively. We think of $H_{t}$ as being the frontier of a regular neighborhood of the spine. Let $S_{0}\subset\mathcal{S}$ be a component. Each intersection of $S_{0}$ with $H$ is either a loop in $C$ parallel to $\partial S_{0}$ or a meridian of an edge of the spine. A meridian intersection bounds an s-disc for $H$ in $C$. If some such disc is a compressing disc, the interval is labelled - (if it is the lowest interval) or ++ (if it is the highest interval). If not, $H_{t}$ is the surface isotopic to $H$ satisfying the conclusions of the lemma.
Assume, therefore, that both sides of $H$ admit compressing discs, so the lowest interval is labelled - and the highest is labelled ++. We conclude that there must be a subinterval with no label. Let $t$ lie in that subinterval. If $H_{t}\cap\mathcal{S}=\varnothing$, we are done, so suppose that $\gamma\subset H_{t}\cap\mathcal{S}$ is a component. As $\mathcal{S}$ is the union of discs and spheres, we may assume that it is innermost in $\mathcal{S}$. Let $D\subset\mathcal{S}$ be the innermost disc it bounds. Since the interval containing $t$ has no label, $\partial D$ is inessential on $H$. If $D$ is not a semi-compressing disc for $H$, we can isotope $H$ so as to remove $\partial D$ from $H\cap\mathcal{S}$. ∎
@@ 4. Heegaard surfaces through thick and thin
Following [HS], we reinterpret [ST] in terms of surfaces, rather than handle structures. Suppose that $\mathcal{H}$ is an oriented generalized Heegaard surface and that $H$ is an weakly reducible thick surface. Let $D_{-}$ and $D_{+}$ be unions of pairwise disjoint s-discs for $H$, with $D_{-}$ below $H$ and $D_{+}$ above $H$. Let $H_{\pm}=H|_{D_{\pm}}$ and let $F=H|_{D_{-}\cup D_{+}}$. A small isotopy of $H_{+}$ above $H$ and $H_{-}$ below $H$ makes $_{-}, H+H_{+}$, and $F$ mutually disjoint. Let $\mathcal{J}^{+}=(\mathcal{H}\setminus H)\cup H_{-}\cup H_{+}$ and let $\mathcal{J}^{-}=\mathcal{H}^{-}\cup F$. Then we say that $\mathcal{J}=\mathcal{J}^{-}\cup\mathcal{J}^{+}$ is obtained by an untelescoping $\mathcal{H}$ and that $D_{-}$ and $D_{+}$ are an s-weak reducing pair. It is a weak reducing pair if both $D_{-}$ and $D_{+}$ contain a compressing disc and a (s/2)-weak reducing pair if at least one of $D_{-}$ or $D_{+}$ contain a compressing disc. See Figure 4.
[…]
@@
On the other hand, suppose $\mathcal{H}$ contains a thick surface $H$ and a thin surface $F$ such that $H\cup F$ lie in the boundary of a product compressionbody component $C$ of $M\setminus\mathcal{H}$. We say that $\mathcal{J}=\mathcal{H}\setminus(H\cup F)$ is obtained by consolidating $\mathcal{H}$.
@@
    untelescoping using an (s/2)-weak reducing pair $_{-}, D+D_{+}$, such that $D_{\pm}$ is a compressing disc and a positive genus component of $H_{\mp}=H|_{D_{\mp}}$ can be consolidated with a component of $\mathcal{H}^{-}$; we also perform that consolidation.
@@
If $\alpha,\beta$ are two such complexity sequences, we say that $\alpha<\beta$ if and only if there exists $k\geq 0$ such that the first $k$ terms of $\alpha$ and $\beta$ coincide and either $\alpha$ does not have a (k+1)(k+1)st term, but $\beta$ does (possibly it is 0) or the (k+1)(k+1)st term of $\beta$ is strictly greater than the (k+1)(k+1)st term of $\alpha$. Observe that complexities are well-ordered.
Consolidation removes a term from a complexity sequence and so decreases it. A type I untelescoping replaces an entry in a complexity sequence with some number (at least two) of other terms, each strictly smaller than the original. Thus, it also decreases complexity. Finally, consider a type II untelescoping and consolidation. Assume the untelescoping uses an (s/2)-weak reducing pair $_{-}, D+D_{+}$ such that $D_{-}$ contains only semi-compressing discs and $D_{+}$ contains at least one compressing disc. Let $H\subset\mathcal{H}^{+}$ contain $\partial D_{-}\cup\partial D_{+}$. Since $H$ admits a compressing disc, $c(H)>0$. Then if $H^{\prime}$ is a component of $H_{+}=H|_{D_{+}}$, we have $c(H^{\prime})<c(H)$. If $H^{\prime}$ is a component of $H_{-}=H|_{D_{-}}$, either $c(H^{\prime})=c(H)$ or $c(H^{\prime})=0$. In the former case, we immediately consolidate the component $H^{\prime}$, and so the type II untelescoping and consolidation also strictly decreases complexity. ∎
@@
Suppose that $\mathcal{H}$ is a locally thin oriented generalized Heegaard surface for $M$. Then the following hold:
@@
If $\mathcal{H}$ is locally thin, we cannot guarantee that each component of $\mathcal{H}$ is essential; there may be components that are 2-spheres bounding 3-balls or components that are $\partial$-parallel. Such thin surfaces will not interfere in what follows. As in [ST, HS, TT1], if we allowed certain destabilization operations, we could continue thinning and ensure the thin surfaces were essential. In this paper, we don’t do this because we want to be able to reconstruct (via amalgamation) a Heegaard surface $H$ from a locally thin generalized Heegaard surface $\mathcal{H}$ to which it thins.
@@
Assume that $M$ is a reducible manifold such that no component of $\partial M$ is a sphere. If $\mathcal{H}$ is a locally thin generalized Heegaard surface for $M$, then $\mathcal{H}^{-}$ contains a collection of essential spheres $P$ such that each component of $M|_{P}$ is irreducible. In particular, if $M$ contains a nonseparating sphere, then so does $\mathcal{H}^{-}$.
@@
Let $P\subset\mathcal{H}^{-}$ be the union of all the spheres in $\mathcal{H}^{-}$ that are essential in $M$. A priori $P$ may be empty, however we will see that it is not. Let $M_{0}$ be a component of $M|_{P}$ and let $\mathcal{H}_{0}$ be those components of $\mathcal{H}$ contained in the interior of $M_{0}$. Using the fact that $\mathcal{H}$ cannot be consolidated and that the existence of spherical components in $\partial_{-}C$ for a compressionbody $C$, does not affect which curves in $\partial_{+}C$ bound compressing discs in $C$, we see that $\mathcal{H}_{0}$ is also thin.
@@
Since the previous paragraph modifies the sphere $Q$ by surgery, and since every sphere in a compressionbody is separating, if $Q$ were nonseparating, we see that $P$ must also have contained a nonseparating sphere. ∎
@@ 5. Amalgamation
Suppose that $M$ is a 3-manifold containing a transversally oriented closed surface $F\subset M$ (possibly disconnected) separating $M$ into (possibly disconnected) 3-manifolds $M_{-}$ and $M_{+}$, such that the transverse orientation on $F$ points into each component of $M_{+}$ and out of each component of $M_{-}$ and so that each component of $M_{-}$ and $M_{+}$ is incident to a component of $F$. Suppose that $H_{\pm}$ is the union of oriented Heegaard surfaces for $M_{\pm}$ such that the orientation on $H_{+}$ (resp. $H_{-}$) points away from (resp. towards) $F$. We can amalgamate the surfaces $H_{-}$ and $H_{+}$ across $F$ as follows. See Figure 5.
[…]
@@
For notational convenience, let $U_{\pm},V_{\pm}$ be the union of the compressionbodies on either side of $H_{\pm}$ with $F=U_{-}\cap V_{+}$. Choose a complete collection of discs for each of $U_{-}$ and $V_{+}$. When we $\partial$-reduce $U_{-}\cup V_{+}$ along those discs, we obtain product compressionbodies and 3-balls. The union of some of those product compressionbodies is $\Pi=F\times[-1,1]$ where $F\times[-1,0]\subset V_{+}$ and $F\times[0,1]\subset U_{-}$. In both cases, consider $F$ as $F\times\{0\}$. The other product compressionbodies contain $\partial_{-}V_{+}\setminus F$ and $\partial_{-}U_{-}\setminus F$. Let $\Delta_{-}\subset U_{-}$ be those discs in the complete collection that leave scars on $F\times\{1\}$. A small isotopy of the product structure on $\Pi$ ensures that the projection of the scars from $\Delta_{-}$ on $F\times\{1\}$ to $F\times\{-1\}$ completely misses the scars in $F\times\{-1\}$ arising from the complete collection of discs in $V_{+}$. Given a complete collection of discs for $V_{-}$, we may also ensure that the projections miss the remnants of their boundaries, as they are arcs in $F\times\{-1\}$. Let $A\subset\Pi$ be the solid tubes corresponding to the projections. Let $X_{-}$ be the union of $V_{-}$ with $A$ and with the components of $U_{-}|_{\Delta_{-}}$ not containing $F$. Let $X_{+}$ be the closure of the complement of $X_{-}$ and let $H=X_{-}\cap X_{+}$. See Figure 5.
@@
Suppose that $M_{-}$ and $M_{+}$ are 3-manifolds such that $F=\partial M_{1}\cap\partial M_{2}$ is nonempty and oriented so that its orientation points into $M_{+}$ and out of $M_{-}$. Let $H_{-}$ and $H_{+}$ be oriented Heegaard surfaces for $M_{-}$ and $M_{+}$ respectively such that the orientation for $H_{+}$ points away from $F$ and the orientation for $H_{-}$ points toward $F$. Let $H$ be the Heegaard surface resulting from amalgamating $H_{+}$ and $H_{-}$ across $F$. Then:
@@
    If $F$ is a sphere, then $H$ is isotopic to the connected sum of $H_{-}$ and $H_{+}$ as Heegaard surfaces for the 3-manifolds obtained from $M_{-}$ and $M_{+}$ by capping off $F$ with 3-balls.
@@
A height function on an oriented generalized Heegaard surface $\mathcal{H}$ is a function $f\mskip 0.5mu\colon\thinspace\mathcal{H}\to\mathbb{N}$ which is constant on each component of $\mathcal{H}$ and satisfies:
@@
Suppose that $\mathcal{H}$ is an oriented acyclic generalized Heegaard surface with a height function $f$. Suppose that $C$ is a compressionbody of $M\setminus\mathcal{H}$ and that $H=\partial_{+}C$. Let
@@
If $C$ is below $H$ (resp. above), let $\mathcal{A}(C)$ be the subset of $\mathcal{C}(C)$ on which $f$ is maximal (resp. minimal), where the optimization is over all thick surfaces in $\mathcal{C}(C)$. Suppose that $\mathcal{J}$ is obtained by an amalgamation of $\mathcal{H}$ and that there exists a compressionbody $C\subset M\setminus\mathcal{H}$ so that $\partial_{+}C$ is amalgamated with all thick surfaces in $\mathcal{A}(C)$, then we say that $\mathcal{J}$ is obtained by an amalgamation of $\mathcal{H}$ that is consistent with the height function $f$.
@@
If $\mathcal{J}$ is obtained from $\mathcal{H}$ by an amalgamation consistent with a height function $f$, then $\mathcal{J}$ is acyclic and $f$ naturally induces a height function for $\mathcal{J}$.
@@
Use the notation of Definition 5.8. Without loss of generality, assume that $C$ is below $H$. This implies that each $C^{\prime}_{i}$ is above $H_{i}$. Let $C_{i}$ be the compressionbody below $H_{i}$ and $C^{\prime}$ the compressionbody above $H$. Let $H^{\prime}$ be the surface resulting from the amalgamation of $H$ and $H_{i}$ across components of $\partial_{-}C$ . Let $U$ and $V$ be the compressionbodies above and below $H^{\prime}$ respectively. Let $\mu=f(H_{i})$ and note this is the same for all $i$.
@@
$f^{\prime}(F)=f(F)>f(H_{i})=f^{\prime}(H^{\prime}).$
@@
[…]
Figure 6. An example of an amalgamation respecting a height function. The numbers indicate the heights of the thick and thin surfaces. The top thick surfaces are amalgamated with the next two thick surfaces in the left two columns but not the next thick surface in the rightmost column. Those are the two thick surfaces that are amalgamated with the topmost thick surface since they have the largest heights of the thick surfaces immediately adjacent to the topmost thick surface. The fact that the top right thin surface has a higher height is irrelevant.
@@
Let $\mathcal{H}$ be an oriented generalized Heegaard surface with height function $f$ and let $s$ be a sequence of amalgamations consistent with the induced height functions. For any compressionbody $C\subset M\setminus\mathcal{H}$ with $C$ below $\partial_{+}C$, if $F$ is a component of $\partial_{-}C$ with $f(F)<f(\partial_{+}C)$, add $(f(\partial_{+}C)-f(F)+1)/2$ parallel copies of $F$ to $\mathcal{H}$. One component in each copy is a thick surface and one is a thin surface and, together with $F$ and $\partial_{+}C$, the thick and thin surfaces alternate. Observe that we can restore $\mathcal{H}$ by consolidating the parallel copies we just inserted. We can extend both the height function $f$ and the sequence of amalgamations to incorporate these new copies. As amalgamating across a product region is the same as consolidation (Lemma 5.2), this does not affect the result. Consequently, the result of amalgamation consistent with a height function, is independent of the height function, depending only on $\mathcal{H}$ and on the order of thick surfaces chosen in $s$. We now show how to eliminate even this latter dependency.
Let $\mathcal{H}^{\prime}$ be the new generalized Heegaard surface after one of the amalgamations. Its dual digraph is the result of merging edges of the dual digraph of $\mathcal{H}$. Observe that if $C$ is a compressionbody below $\partial_{+}(C)$, then
@@
We may as well assume, therefore, that for each term of the amalgamation sequence $s$, there is an $i\in\mathbb{N}$ such that the thick surfaces at height $i$ are amalgamated with the thick surfaces at height i1i-1. It then follows from [Lackenby, Prop. 3.1] that $H$ is, up to isotopy, independent of $s$ and of the choice of disc sets used in each amalgamation. ∎
@@
The case when $\partial M$ contains spheres follows relatively easily from the case when it does not. (Capping off spherical components of $\partial M$ with 3-balls does not affect the Heegaard genus of $M$. It can, however convert an essential sphere $S$ into an inessential sphere. However, if that occurs, the sphere $S$ bounds a punctured 3-ball in $M$, which has Heegaard genus 0.) For simplicity, assume that $\partial M$ does not contain spheres. For a generalized Heegaard surface $\mathcal{H}$, let $\operatorname{net}\chi(\mathcal{H})=-\chi(\mathcal{H}^{+})+\chi(\mathcal{H}^{-})$. It is easy to verify that if $\mathcal{H}$ thins to $\mathcal{J}$ or if $\mathcal{J}$ amalgamates to $\mathcal{H}$, then $\operatorname{net}\chi(\mathcal{H})=\operatorname{net}\chi(\mathcal{J})$. We treat the case of separating and nonseparating spheres separately; in each case, proving two inequalities.
@@
Suppose that $M$ contains a nonseparating sphere $P$. Let $H$ be a minimal genus Heegaard surface for $M|_{P}$. The scars from $P$ are two 3-balls $B_{1},B_{2}$ in $M|_{P}$. We may isotope $H$ so that those 3-balls lie on the same side of $H$. On that side of $H$, choose a 3-ball $B$ containing those scars. Remove the interiors of $B_{1},B_{2}$ creating two spherical boundary components $F_{1},F_{2}$ of $M\setminus P$. Let $W=S^{2}\times I$ and let $H^{\prime}\subset W$ be a sphere bounding a 3-ball in $W$. Let $M^{\prime}$ be the result of gluing the components of $\partial W$ to $F_{1}\cup F_{2}$. Then it is possible to assign orientations to $\mathcal{H}=H\cup F_{1}\cup F_{2}\cup H^{\prime}$ so that it is an oriented, acyclic generalized Heegaard surface with $H,H^{\prime}\subset\mathcal{H}^{+}$ and $F_{1},F_{2}\subset\mathcal{H}^{-}$. Notice that $M^{\prime}$ is homeomorphic to $M$ and that